# Adaptive Blind All-in-One Image Restoration

### v1

@@ 4 Experiments
We evaluate our method on two setups for all-in-one IR: five- and three-degradation setups. Additionally, we test on datasets excluded from training, novel IR degradation types, and mixed degradation scenarios. The accuracy of the methods is assessed using two well-established metrics: PSNR and SSIM. In all cases we report the mean value for all the test images, and we highlight best and second-best values for each metric. Our approach is compared against recent all-in-one IR methods. Notably, some current methods are either evaluated under different setups or lack available code and models, which complicates direct comparisons. Therefore, we train three state-of-the-art methods, Restormer [63], PromptIR [40], and X-Restormer [8], on the five-degradation IR setup to provide a fair comparison. We used the authors’ provided code. For our approach, we evaluate the two variations based on the variants of the estimator, namely one-hot (OH) and soft weights (SW).
@@
Our training is conducted in three phases using Adam [23] with weight decay [29], an initial learning rate of $2\times 10^{-4}$, a cosine learning rate scheduler, and a warmup start of one epoch. In phase I, we apply a weighted combination of cross-entropy loss $L_{CE}$ for per-pixel degradation maps and a reconstruction loss $L_{R}$, where $L_{R}$ consists of $L_{1}$ loss and SSIM loss. Both the cross-entropy and SSIM components are weighted with 0.5 In phase II, we use only the reconstruction loss, where, as before, SSIM loss is weighted with 0.5. In phase III, we apply cross-entropy loss between the regressor’s output and the input image’s degradation type.
Table 4: Quantitative results for unseen IR tasks. Note that the models have not been trained for these degradations. Ours* shows results for the lightweight re-training scenario. New adapters are trained for the new tasks and the estimator is retrained with 8 tasks (5-IR case + 3 new ones; only 8M training parameters).
@@
5-Degradation blind IR:
For the blind five-task setup, we follow the protocol by Zhang et al. [64]. Specifically, we use Rain200L[56] for deraining, RESIDE [25] for dehazing, BSD400 [33] and WED [32] for denoising with $\sigma=25$, GoPro [35] for deblurring, and LOL [52] for low-light enhancement. For evaluation, we employ Rain100L [56], SOTS-Outdoor [25], BSD68 [33], GoPro [35], and LOL [52]. Results are reported in Tab. 1. Our approach outperforms all state-of-the-art methods across all tasks, with the exception of PSNR for denoising. Notably, our method achieves a substantial improvement over the state of the art, yielding a 2.91 dB average PSNR gain across the five tasks. In particular, our method surpasses IDR [64] by 2.16 dB on image deraining and provides a 2.40 dB improvement over PromptIR [40] on low-light image enhancement.
GoPro [35] Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
LoLv1 [52] Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Input Restormer [63] PromptIR [40] Ours-OH Ground truth
@@
Following Li et al. [27], we further evaluate our approach on a 3-task blind IR setup, comparing our approach with specialized all-in-one methods for deraining, dehazing, and denoising. Compared to the five-task setup, we omit deblurring and low-light enhancement and instead introduce two additional noise levels: $\sigma=15$ and $\sigma=50$. Two LoRA adapters for these noise levels are trained, and the estimator is retrained for the new setting. Results for this 3-task setup are presented in Tab. 2. Our approach outperforms all the other state-of-the-art methods on average and stays consistent in all the degradation types.
JPEG [42] Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
4-to-8 bits [42] Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Input PromptIR [40] Ours-SW Ours-SW retrained Ground truth
@@
Figure 5: Qualitative results for unseen IR tasks, including JPEG artifact removal and 4-to-8 bit reconstruction. PromptIR [40] and Ours are not trained for this task, while Ours retrained has a specified LoRA in an 8-degradation setup.
Additional test sets:
@@
To assess the generalization capabilities of our approach, we evaluate the model on three datasets not used during training. Specifically, we use the Rain100H [56] dataset for heavy deraining, HIDE [43] for human-centric deblurring, and LoLv2-Real [58] for low-light image enhancement. The results of the three best methods on the 5-degradation setup and our approach are presented in Tab. 3. Our method outperforms all other approaches across all datasets and metrics. Notably, we achieve a significant improvement of 7.19 dB over Restormer [63] on Rain100H [56]. This performance boost can be attributed to the pre-training on synthetic data, which allows our model to learn from a broader set of natural images and degradation scenarios compared to the other methods’ small training set, like Rain200L [56]. Additionally, we achieve over 1 dB PSNR improvement on both HIDE and LoLv2-Real, further demonstrating the effectiveness and robustness of our approach.
Unseen IR tasks:
Although our approach was initially trained on five types of degradations, we further evaluate its generalization to three additional IR tasks not encountered during training. Specifically, we assess our method on JPEG artifact removal and bit-depth reconstruction using the Live1 dataset [42], and on desnowing using CityScapes-Snow-Medium [65]. Results for the top four previous state-of-the-art methods from the five-degradation setup, alongside our approach, are presented in Tab. 4. Remarkably, our method outperforms all other approaches across all tasks and metrics. To show the adaptability of our approach, we additionally train a separate adapter for each new task, along with a new estimator that blends all eight task-specific adapters, thereby expanding the model’s versatility. The last two rows of Tab. 4 report results for the newly integrated tasks — shown with a . Notably, with the addition of these task-specific adapters, our model achieves a notable boost in performance, while maintaining an average PSNR of 30.08 dB across the original five IR tasks, outperforming the other state-of-the-art methods trained on the initial setup.
Blur&Noise Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Haze&Snow Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Input Restormer [63] PromptIR [40] Ours-SW Ground truth
@@
Mixed degradations:
The limited availability of datasets with mixed degradations presents a challenge for training IR models under these complex conditions. To evaluate our approach, we evaluate it on three mixed-degradation scenarios: blur combined with noise $\sigma{=}25$ using the GoPro dataset[35], blur with JPEG artifacts using the REDS dataset [36], and haze combined with snow using the SRRS dataset [6]. Notably, some degradations, such as JPEG artifacts and snow, are unseen during training, further increasing the difficulty of these tests. As shown in Tab. 5, our model consistently outperforms other methods across all scenarios and metrics. Handling blur with noise is particularly challenging due to their contrasting nature: blur affects low frequencies, while noise affects high frequencies. Even though, our model achieved around 2dB improvement while others remained at the level of 22dB.
@@
Table 6: Ablation studies on types of pre-training for Phase I, and the rank of LoRA [19] for Phase II and III.
Pre-training PSNR SSIM
IR datasets 28.50 .892
GLD+synth. 30.63 .913
+ CutMix 31.09 .920
+ Aux. segm. 31.17 .921
Rank PSNR SSIM Params
4 31.17 .921 3.6M
8 31.14 .920 7.2M
16 30.97 .916 14.3M
Qualitative results:
In Fig. 4, we present examples featuring three types of single degradations from the five-task setup on which the model was trained. Our approach effectively enhances the license plate in the first row, reduces noise in low-light conditions in the second row, and removes heavy rain streaks in the third. In Fig. 5, we show unseen IR tasks including JPEG artifact removal and 4-to-8-bit reconstruction using the Live1 dataset [42]. Our method successfully removes JPEG artifacts from the parrot’s plumage and artifacts caused by bit-depth reduction from the cloud and blue sky. By training a new adapter for these tasks (ours-SW retrained), we achieve superior results with minimal additional training time. Finally, in Fig. 6, we demonstrate our method’s performance under mixed degradations. The first row shows that our method achieves the best reconstruction of the text in the image, while the second row illustrates our model’s effectiveness in removing both haze and snowflakes.
Ablation studies:
In Tab. 6, we perform an ablation study to assess the contribution to the final performance of the proposed pre-training pipeline and the LoRAs’ rank. For pre-training in Phase I, we use the same setup as the five-task configuration, i.e. training the LoRAs and the estimator using task-specific IR datasets and the obtained pre-trained weights. The first row — IR datasets — reflects pre-training on standard IR benchmarks, yielding performance comparable to state-of-the-art methods — see Tab. 1. In contrast, simply pre-training on GLD with synthetic degradations improves performance by 2.15 dB PSNR, demonstrating the importance of Phase I to provide the model with strong generalization capabilities. Further incorporating both the CutMix-like [61] strategy and the auxiliar segmentation task provides an additional performance boost. We also compare LoRA [19] with different ranks for Phase II and III. The model achieves the best performance and efficiency when the rank is set to 4, which we adopted as our default setting. For more ablation results, please refer to our Supplementary Material.
@@ 5 Conclusion
In this work, we introduced an adaptive blind all-in-one IR model aiming towards practical IR. We design both dedicated adapters for strongly handling specific distortions but also a flexible architecture for dealing with practical IR challenges.We first developed a pre-training pipeline featuring multiple synthetic degradations over a large dataset that boosts the generalization of the model. Second, we derived compact per-task adapters that robustly adapt to specific degradations. Third, we developed a lightweight degradation estimator that identifies varying degradations to blend the respective adapters, which is also capable of efficiently incorporate new degradations by training a small fraction of parameters. Our model largely outperforms the state-of-the-art on five- and three-task IR setups, and shows improved generalization to unseen datasets and IR tasks.
@@ Appendix B Synthetic Degradations Generation
As with rainy images, capturing paired degraded and clean images under hazy conditions is nearly impossible due to the variability of environmental factors. Inspired by prior works [26, 3, 16], we model haze degradation using the Kochsmieder model [17], which describes how the visibility of distant objects diminishes, vanishing into the horizon as a function of their distance from the observer. We can formulate it as:
@@
where $I_{input}$ is the input clean image, $T$ is the transmission map derived from the estimated depth map, $A$ is the airlight, and $I_{haze}$ is the resulting image with synthetic haze.
We estimate the depth map of the input image using DepthAnythingv2 [55]. This depth map is normalized within the predefined minimum and maximum haze values to generate the transmission map $T$. The depth map is then replicated for each color channel to compute Eq. 3. To ensure variability, all parameters are randomly sampled for each forward pass. The parameter ranges are as follows: minimum haze $[0.2,0.4]$, and maximum haze $[0.7,0.9]$, airlight is unique for the three color channels with values ranging $[140,200]$.
@@
Low-Light:
When an image is captured under low-light conditions, cameras amplify the sensor’s signal to capture details, amplifying noise, and leading to grainy or speckled artifacts. Furthermore, the reduced dynamic range limits the ability to capture the range of intensities, resulting in color inconsistencies, and loss of detail in shadows and highlights. To simulate these conditions, we compress the input image histogram by a factor $c$ and add a noise with a meager $\sigma$ value. This process can be expressed mathematically as:
$I_{lol}=I_{input}\cdot c+\mathcal{N}(0,\sigma).$ (6)
@@ Appendix C Additional Phase II Ablation Studies
Ours OH 37.73 .978 33.46 .983 31.38 .898 29.00 .878 24.20 .865 31.15 .920
Ours SW 37.79 .979 33.48 .984 31.38 .898 29.00 .878 24.19 .865 31.17 .921
@@ Appendix D Additional Phase III Analysis
In the third phase of our method, we train an estimator to select or blend the most suitable set of adapters based on the input image. The estimator architecture comprises four blocks of Conv2D layers, each followed by batch normalization, ReLU activation, and max pooling, culminating with a global average pooling layer and a linear projection. The total parameter count for the estimator is 538K.
@@
Finally, to highlight the importance of our estimator, we perform a last ablation comparing our approach to just adding or averaging the five degradation task-specific adapters. Tab. 8 shows the results of this ablation. When all adapters are added, the $\Delta W$ values become excessively large for the baseline model to effectively handle them. On the other hand, averaging the adapters results in suboptimal restored images, as the model fails to specialize in any particular degradation type.

### v2

@@ 4 Experiments
We evaluate our method on two all-in-one IR setups: five- and three-degradation. Additionally, we assess its performance on datasets excluded from training, unseen IR degradation types, and mixed degradation scenarios. The accuracy of the methods is measured using two well-established metrics: PSNR and SSIM. In all cases we report the mean value across all the test images, and we highlight best and second-best values for each metric. Our approach is compared against recent all-in-one IR methods. However, some existing methods have been evaluated under different setups or lack publicly available training code and models [25, 2, 62]. To ensure a fair comparison, we train five state-of-the-art methods, Restormer [76], DA-CLIP [37], DiffUIR [38], PromptIR [49], X-Restormer [8], on the five-degradation IR setup. We used the official code provided by the authors. For our approach, we evaluate the two variations, one-hot (OH) and soft weights (SW).
@@
Our training follows a three-phase scheme, using the Adam [28] optimizer with weight decay [36], an initial learning rate of $2\times 10^{-4}$, a cosine learning rate scheduler, and a warmup start of one epoch. Following Potlapalli et al. [49], we train our model on four NVIDIA A40 GPUs with a batch size of 4 and image patches of 384. In the Phase I, we jointly optimize a weighted combination of two objectives: a cross-entropy loss $L_{CE}$ for per-pixel degradation classification and a reconstruction loss $L_{R}$, which consists of $L_{1}$ loss and SSIM loss. The cross-entropy and SSIM components are both weighted at 0.5 to balance classification and reconstruction. During Phase II, we optimize only the reconstruction loss $L_{R}$, maintaining the same SSIM weighting as in the previous phase. In Phase III, we train the lightweight estimator separately using a cross-entropy loss to classify the degradation type of the input image, ensuring that the model accurately learns degradation prediction.
Table 4: Quantitative results for unseen IR tasks. Note that the models have not been trained for these degradations. Ours* shows results for the lightweight re-training scenario. New adapters are trained for the new tasks and the estimator is retrained with 8 tasks.
@@
5-Degradation blind IR.
For the blind five-task setup, we follow the protocol of Zhang et al. [78]. Specifically, for training we use Rain200L[69] for deraining, RESIDE [30] for dehazing, BSD400 [41] and WED [40] for denoising with $\sigma{=}25$, GoPro [43] for deblurring, and LOL [64] for low-light enhancement. Evaluation is conducted on Rain100L [69], SOTS-Outdoor [30], BSD68 [41], GoPro [43], and LOL [64], with results reported in Table 1. We outperform all the methods across all tasks, except PSNR in denoising. Note that denoising gains are typically small due to the identical noise distribution in training and test sets in all-in-one setups. Our method achieves a substantial improvement over the state of the art average, with a PSNR gain of 1.05 dB. Remarkably, our method outperforms AdaIR [10] by 2.95 dB PSNR and 1.19 dB PSNR on dehazing and low-light image enhancement, respectively.
GoPro [43] Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
LoLv1 [64] Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Rain100H [69] Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Input Restormer [76] PromptIR [49] AdaIR [10] Ours-OH Ground truth
@@
Following Li et al. [32], we evaluate our approach on a three-task blind IR setup, comparing it to specialized all-in-one methods for deraining, dehazing, and denoising. Compared to the five-task setup, we omit deblurring and low-light enhancement while introducing two additional noise levels: $\sigma{=}15$ and $\sigma{=}50$ on the BSD68 [41] dataset. Two LoRA adapters and the lightweight estimator are trained for these new noise levels. The baseline model remains the same, trained on the five-task setup. Results in Table 2 demonstrate that our approach outperforms all state-of-the-art methods on average while maintaining consistent performance across all degradation types. In detail, we achieve 0.52 dB average PSNR over AdaIR [10], even when our baseline was trained on the five-degradation setup.
JPEG [53] Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
4-to-8 bits [53] Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Input PromptIR [49] AdaIR [10] Ours-SW Ours-SW retrained Ground truth
@@
Figure 5: Qualitative results for unseen IR tasks, including JPEG artifact removal and 4-to-8 bit reconstruction. PromptIR [49], AdaIR [10], and Ours are not trained for these tasks, while Ours-retrained has a specified LoRA in an 8-degradation setup.
Additional test sets.
@@
To assess generalization, we evaluate our model on three datasets not used during training: Rain100H [69] for heavy deraining, HIDE [54] for human-centric deblurring, and LoLv2-Real [71] for low-light image enhancement. Table 3 presents the results for the methods trained in the five-task setup. Our method outperforms all competitors across all datasets and metrics, achieving substantial improvements of 3.98 dB over DiffUIR [81] on Rain100H, 1.65 dB over X-Restormer [8] on HIDE, and 1.39 dB over PromptIR [49] on LoLv2-Real. These results highlight the effectiveness and robustness of our approach, which benefits from pretraining on synthetic degradations, enabling it to generalize across diverse natural images and degradation scenarios.
Unseen IR tasks.
Although our model is trained on five degradation types, we assess its adaptability to three additional IR tasks: JPEG artifact removal and bit-depth reconstruction using the Live1 dataset [53], and desnowing using CityScapes-Snow-Medium [79]. The results in Table 4 demonstrate that our method consistently outperforms competing approaches. Moreover, by leveraging its modular scheme, we train independent LoRA adapters for each new task, blending them with the five-task baseline (denoted with in Table 4). This strategy allows for refining the model with minimal training (3.6M parameters per task) while preserving previously acquired knowledge, enabled by our robust baseline model and three-phase training procedure. In contrast, other methods either require retraining the entire model from scratch or fine-tuning it without ensuring that prior knowledge is retained.
Blur & N Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Haze & Snow Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Input Restormer [76] PromptIR [49] AdaIR [10] Ours-SW Ground truth
@@
Mixed degradations.
We also evaluate our approach on three mixed-degradation scenarios: blur with noise ($\sigma{=}25$) using GoPro [43], blur with JPEG artifacts using REDS [44], and haze with snow using SRRS [5]. Notably, JPEG artifacts and snow are unseen during training. As shown in Table 5, our model outperforms all competitors by at least 1 dB PSNR in all three cases. Addressing blur and noise simultaneously is particularly challenging due to their opposing frequency characteristics. Our specialize-then-merge scheme allows the lightweight estimator to efficiently blend the task-specific adaptors, offering greater flexibility than end-to-end models trained on fixed degradation types.
Qualitative results.
In Figure 4, we present examples featuring three types of single degradations from the five-task setup. Our approach effectively enhances the license plate in the first row, reduces noise in low-light conditions in the second row, and removes heavy rain streaks in the third. In Figure 5, we show unseen IR tasks including JPEG artifact removal and 4-to-8-bit reconstruction using the Live1 dataset [53]. Our method successfully removes JPEG artifacts from the parrot’s plumage and artifacts caused by bit-depth reduction from the cloud and blue sky. By training a new adapter for these tasks (ours-SW retrained), we achieve superior results with minimal additional training. Finally, in Figure 6, we demonstrate our method’s performance under mixed degradations. The first row shows that our method achieves the best reconstruction of the text in the image, while the second row illustrates our model’s effectiveness in removing haze and snowflakes.
@@
Table 6: Ablation studies on pretraining strategies for our method (a), and PromptIR [49] (b). We report the five-task setup average.
(a) ABAIR PSNR SSIM
IR datasets 29.50 .892
GLD+2 global 27.68 .887
GLD+1 global 30.63 .913
GLD+CutMix 31.11 .920
+ Aux. seg. 31.21 .921
(b) PSNR SSIM
PromptIR [49] 28.58 .885
RAM [50] 28.79 .889
Art [66] 28.83 .889
GLD+1 global 29.27 .902
GLD+CutMix 29.58 .908
Ablation studies.
Our method consistently outperforms previous state-of-the-art approaches across seen, unseen, and mixed degradations, as demonstrated in all the tables. To further assess its effectiveness, we conduct additional experiments regarding the impact of our Phase I pretraining strategy. Table 6 reports the results averaged over the five-task setup.
First, we evaluate the impact of the synthetic pretraining to the final performance of our method. Table 6 (a) compares different pretraining setups: standard IR datasets as pretraining (IR datasets), Google Landmarks dataset with one and two simultaneous global synthetic degradations applied across the entire image, (GLD+1 global and GLD+2 global), respectively, as well as incorporating the Degradation CutMix (GLD+Cutmix) and the auxiliary segmentation head with its corresponding loss (Aux. seg.). Next, we compare our synthetic pretraining approach with the pretraining strategy from RAM [50] and the optimization technique from Art [66] on the PromptIR [49] model. We select PromptIR as a well-established baseline for evaluation. The results of these comparisons are presented in Table 6 (b).
From Table 6, we derive four key observations: (i) Our synthetic pretraining substantially outperforms other pretraining and optimization techniques. (ii) Our three-phase approach is effective, as even without using synthetic pretraining (IR datasets), our method surpasses the baseline PromptIR model. (iii) Training with multiple simultaneous degradations (GLD+2 global) negatively impacts performance, likely due to excessive degradation in Phase I, which hinders the learning of robust baseline. (iv) Both Degradation CutMix and the auxiliary segmentation head contribute substantially to performance improvements. These findings highlight the importance of both synthetic pretraining and the three-phase training scheme in achieving optimal results. Additional results and ablation studies for Phases II and III, including analyses of different adapter types, ranks, and alternative strategies for blending the adapters, are provided in the Supplementary Material.
@@ 5 Conclusion
In this work, we introduce an adaptive blind all-in-one IR (ABAIR) model aiming towards practical IR. We design a specialize-then-merge scheme with dedicated adapters for strongly handling specific distortions and a flexible architecture for dealing with unseen and mixed degradations. We first developed a pretraining pipeline featuring multiple synthetic degradations over a large dataset that boosts the model’s generalization. Second, we derived compact per-task adapters that robustly adapt to specific degradations. Third, we developed a lightweight degradation estimator that identifies varying degradations to blend the respective adapters. Our method is also capable of efficiently incorporate new degradations by training a small fraction of parameters. Our model largely outperforms the state-of-the-art on five- and three-task IR setups, and shows improved generalization to unseen datasets and IR tasks.
@@ Appendix B Synthetic Degradations Generation
As with rainy images, capturing paired degraded and clean images under hazy conditions is nearly impossible due to the variability of environmental factors. Inspired by prior works [31, 17], we model haze degradation using the Kochsmieder model [21], which describes how the visibility of distant objects diminishes, vanishing into the horizon as a function of their distance from the observer. We can formulate it as:
@@
where $I_{input}$ is the input clean image, $T$ is the transmission map derived from the estimated depth map, $A$ is the haze color, and $I_{haze}$ is the resulting image with synthetic haze.
We estimate the depth map of the input image using DepthAnythingv2 [68]. This transmission map is normalized within the predefined minimum and maximum haze values to generate the transmission map $T$. The depth map is then replicated for each color channel to compute Eq. 3. To ensure variability, all parameters are randomly sampled for each forward pass. The parameter ranges are as follows: minimum haze $[0.2,0.4]$, and maximum haze $[0.7,0.9]$, haze color is unique for the three color channels with values ranging $[140,200]$.
@@
Ours OH 38.18 .983 33.46 .983 31.38 .898 29.00 .878 24.20 .865 31.24 .921
Ours SW 38.22 .984 33.48 .984 31.38 .898 29.00 .878 24.19 .865 31.25 .922
Ours Oracle 39.09 .981 33.54 .984 31.40 .901 29.10 .879 24.45 .866 31.39 .913
Low-Light:
When an image is captured under low-light conditions, cameras amplify the sensor’s signal to capture details, amplifying noise, and leading to grainy or speckled artifacts. Furthermore, the reduced dynamic range limits the ability to capture the range of intensities, resulting in color inconsistencies, and loss of detail in shadows and highlights. To simulate these conditions, we compress the input image histogram by a factor $c$ and add a noise with a small $\sigma$ value. This process can be expressed mathematically as:
$I_{lol}=I_{input}\cdot c+\mathcal{N}(0,\sigma).$ (6)
Appendix C Additional Phase II Ablation Studies
@@ Appendix D Additional Phase III Analysis
In the third phase of our method, we train an estimator to integrate the set of adapters with the baseline model based on the input image. The estimator architecture comprises four blocks of Conv2D layers, each followed by batch normalization, ReLU activation, and max pooling, culminating with a global average pooling layer and a linear projection. The total parameter count for the estimator is 538K.
@@
To highlight the importance of our estimator, we perform another ablation comparing our approach to just adding or averaging the five degradation task-specific adapters. Tab. 8 shows the results of this ablation. When all adapters are added, the $\Delta W$ values become excessively large for the baseline model to effectively handle them. On the other hand, averaging the adapters results in suboptimal restored images, as the model fails to specialize in any particular degradation type. Finally, Table 8 also presents the results for the five-task setup assuming the lightweight estimator is perfect, i.e., an oracle. The results are only slightly better from those obtained when using our estimator, confirming that its performance is strong. Please note that these Oracle results can be interpreted as single-task performance, as the selected adapter always corresponds to the specific degradation type of the input image.